The area of each plate of a parallel plate capacitor is 0.03m2. The distance between the plates of the capacitor is 2×10-3m, and the potential difference between the plates is 150V. Calculate capacitance of the capacitor.
1
Expert's answer
2020-06-03T12:05:47-0400

Solution

"C=\frac{\epsilon\times \epsilon_0\times S}{4\times\pi\times d};\\dielectric\; permittivity\; of\; air\;\epsilon=1,\;so\\C=\frac{ \epsilon_0\times S}{4\times\pi\times d};\\C=\frac{ 8.85\times10^{-12}\times 0.03}{4\times\ 3.14\times 2\times10^{-3}};\\C=10.57\;pF"

Answer: capacitance of the capacitor equals 10.57 pF